OpenStudy (annie12m):

Medal to check my work: Melinda starts with $22 and saves $10 each week. George starts with $10 and saves $13 each week. After how many weeks will they have the same amount saved? 22 + (10 * x) = 10 + (13 * x) -10 -10 =12 12 + (10 * x) = 13 * x -12 -12 =1 10x = 1x /1 /1 10 = x? Correct? Probably not, will someone help me?

OpenStudy (annie12m):

22 + 10 * 10 = 10 + 13 * 10 100 130 22 + 100 = 10 + 130 122 = 140 not correct

OpenStudy (anonymous):

This part is not right 12 + (10 * x) = 13 * x -12 -12 =1 Can't subtract 12 from the coefficient of x

OpenStudy (b87lar):

in the second step you should subtract 10x , instead of 12

OpenStudy (annie12m):

12 + 10x = 13x /10 /10 =1.3 12 = 1.3x /1.3 /1.3 9.23076923077 = x? or are you suppose to subtract 1.3? 12 = 1.3x -1.3 -1.3 10.7 = x???

OpenStudy (annie12m):

wait your suppose to subtract 10?

OpenStudy (annie12m):

12 +10x = 13x -10 -10 3x 12 = 3x 3 3 4 = x

OpenStudy (annie12m):

@aleroth @b87lar

OpenStudy (anonymous):

That's not right either, you want to get all numbers with x on one side and all 'lonely' numbers on the other side

OpenStudy (anonymous):

What b87lar meant was to subtract 10x on both sides

OpenStudy (annie12m):

i did that

OpenStudy (anonymous):

Oops, sorry I didn't see the last thing there. Yeah, that's right!! D:

OpenStudy (annie12m):

ty!!